Be sure to answer all parts.

How would you prepare 315 mL of a 0.886 M NaOH solution, starting with a 4.52 M stock solution?

Volume of stock solution needed: ?mL

Volume of water needed: ?mL
Answer #1

Molarity = No of moles/ Volume in L

0.886 M NaOH = 0.886 moles of NaOH in 1L ter

Let,

M1 = 0.886M

V1 =315mL= 0.315 L

M2 = 4.52 M

When a concentrated solution is diluted, the volume increases and the no of moles remain the same. Thus,

M1*V1 = M2*V2

0.886 * 0.315 = 4.52*V2

V2 = 0.0617 L

Thus, to make 315 mL of solution, 61.7 i.e. 62 mL of stock solution will be used while 253 mL of water will be added.